# Network inference via approximate Bayesian computation. Illustration on a stochastic multi-population neural mass model

### v3

@@ 1 Introduction
When embedded into a statistical inference procedure, this standard numerical scheme may therefore yield wrong estimation results, make the inference algorithm computationally infeasible or lead to ill-conditioned estimation methods [7, 12, 28, 30]. To obtain a reliable and efficient numerical simulation method, we further develop the structure-preserving splitting procedure proposed in [1] to our $N$-population SDE model (second contribution). In contrast to the Euler-Maruyama scheme, which is based on truncating a stochastic Taylor series, the idea behind the splitting approach is to divide the unsolvable equation into solvable subequations, and to compose their solutions in a proper way [5, 24]. The constructed splitting scheme for the stochastic $N$-population neural mass model is based on a Hamiltonian type re-formulation of the SDE and successfully handles multiple interacting components, allowing to accurately simulate the complex dynamics of coupled neural populations.
@@
[…]
@@ 2.1 One neural population
The meaning and typical values of the model parameters $A$, $B$, $a$, $b$, $C_{1}$, $C_{2}$, $C_{3}$, $C_{4}$, $\nu_{\max}$, $\gamma$ and $v_{0}$ are reported in Table 1 (see [1, 18]). The average incoming membrane potential of the main cells, i.e., the process
$y(t):=x_{2}(t)-x_{3}(t),$
@@
Table 1: Standard parameter values for the Jansen and Rit Neural Mass Model.
Parameter Meaning Standard value
$A$ Average excitatory synaptic gain $3.25$ mV
@@
$C_{1}$, $C_{2}$ Avg. no. of synaptic contacts in the excitatory feedback loop $C$, $0.8\ C$
$C_{3}$, $C_{4}$ Avg. no. of synaptic contacts in the inhibitory feedback loop $0.25\ C$, $0.25\ C$
$\nu_{\max}$ Maximum firing rate (Maximum of the sigmoid function) $5$ s-1
$v_{0}$ Value for which $50\%$ of the maximum firing rate is attained $6$ mV
@@
An illustration of a simulated trace with parameters chosen to produce $\alpha$-waves (neural oscillations in the 8–12 Hz frequency band) is provided in Figure 1A. The model can also produce more complex behavior, such as brain signals occurring before and during epileptic seizures, by increasing the excitation-inhibition-ratio $A/B$ [42]. In Figure 1B, regular EEG activity is produced by using $\mu=90$, $\sigma=500$, $A=3.25$ and the standard values reported in Table 1. Increasing $A$ causes the model to generate sporadic and frequently occurring spikes ($A=3.5$ and $A=3.6$) and rhythmic discharge of spikes ($A=4.3$), see Figure 1C-1E.
Figure 1: Traces of one neural population with different types of activity. Simulated paths of the process $Y$ (4) of model (3). Panel A: Trace showing $\alpha$-rhythmic activity using the standard values of Table 1, $A=3.25$, $C=134.263$, $\mu=202.547$, and $\sigma=1859.211$ (values taken from [7]). Panels B-E: Traces describing activity occurring during epileptic seizures using the standard values of Table 1, $\mu=90$, $\sigma=500$ and different values of $A$. Panel B: Regular EEG, $A=3.25$. Panel C: Sporadic spikes, $A=3.5$. Panel D: Frequently occurring spikes, $A=3.6$. Panel E: Rhythmic discharge of spikes, $A=4.3$. The signals resemble experimental stereo EEG recordings (cf. Figure 3 in [42]).
@@ 2.2 Multiple coupled neural populations
$dX^{k}_{5}(t)=\Bigl{[}A_{k}a_{k}\Bigl{(}\mu_{k}+C_{2,k}\textrm{sig}\left(C_{1,k}X^{k}_{1}(t)\right)+\sum\limits_{j=1,j\neq k}^{N}\rho_{jk}K_{jk}X_{1}^{j}(t)\Bigr{)}-2a_{k}X^{k}_{5}(t)-a_{k}^{2}X^{k}_{2}(t)\Bigr{]}dt+\sigma_{k}dW^{k}_{5}(t).$ (5)
@@
where $L>0$ is a coupling strength parameter and the parameter $10\ll c<$ 1 describes how fast the network coupling strength decreases with increasing distance between populations.
@@
The parameters of the stochastic multi-populatuion JR-NMM thus consist of continuous model parameters (cf. Table 1), two further continuous coupling strength parameters $L>0$ and $10\ll c<$ 1 (cf. (5) and (6)) and binary coupling direction (network) parameters $\rho_{jk}$ $j,k=1,\ldots,N$, $j\neq k$ (cf. (5)). Moreover, the observed component of the model is given by the $N$-dimensional process
@@
Similar to Section 2.1, we present a short simulation study where we generate paths of $Y$ (7), focusing on the newly introduced coupling term. According to [42], not only the excitation-inhibition-ratio $A/B$ is relevant for epileptic behavior, but also the coupling strengths and directions between neural groups play a crucial role. This is illustrated in Figure 2, where simulated activity of $N=4$ neural populations under different coupling regimes is shown. In the left panels, no coupling occurs, i.e., all $\rho_{jk}$-parameters are set to zero. In the middle and right panels, there is a unidirectional (cascade) coupling structure (illustrated in Figure 3a of Section 4), i.e, $\rho_{12}=\rho_{23}=\rho_{34}=1$, for different coupling strengths $L=K_{12}=K_{23}=K_{34}$. The activity of a passive site (no epileptic spikes occur without input from other populations) strongly depends on that of an active site (epileptic spikes occur without input from other populations). In Population $1$, we set $A_{1}=3.6$ to obtain spiking activity. In the remaining populations, the standard values of Table 1, $\mu=90$ and $\sigma=500$ are used. When there is no coupling, no activation of Populations 2–4 occurs (left panels). Introducing coupling and setting $L=K_{12}=K_{23}=K_{34}=300$ (central panels) leads to a dependence of Populations 2–4 on Population $1$. When the coupling is strong enough ($L=K_{12}=K_{23}=K_{34}=500$, right panels), rhythmic synchronization occurs. A similar behavior for two populations has also been observed for the original JR-NMM (cf. Figure 6 in [42]).
Figure 2: Cascade network of four populations with one active population. Simulated paths of $Y$ (7) for $N=4$ neural populations. The standard values of Table 1, $\mu=90$ and $\sigma=500$ are used, except for Population $1$ where $A_{1}=3.6$ to activate it. Panel A: $\rho_{jk}=0$, $j,k=1,\ldots,4$, $j\neq k$. Panels B-C: $\rho_{12}=\rho_{23}=\rho_{34}=1$. The coupling strength parameters $K_{12}$, $K_{23}$ and $K_{34}$ equal $L=300$ and $L=500$ in Panel B and Panel C, respectively.
@@ 3.1 Standard ABC schemes
[…]
@@
[…]
@@
This procedure is then repeated over several iterations (with decreasing threshold $\delta_{r}<\delta_{r-1}$ chosen as a certain percentile of the distances computed at the previous iteration $r-1$) until a suitable stopping criterion is reached (e.g., the acceptance rate of particles drops below a prefixed level or a certain number of model simulations, denoted by $n_{\textrm{sim}}$, has been reached). The final SMC-ABC posterior is obtained by sampling from the particles kept at the last iteration $r_{\textrm{last}}$ with probabilities given by the corresponding normalized weights. A detailed formulation of the standard SMC-ABC can be found, e.g., in Algorithm 2 of [27].
@@ 3.2 Adapted SMC-ABC algorithm for network inference: nSMC-ABC
At iteration $r=1$, we run the standard acceptance-rejection ABC, sampling parameter candidates $\theta_{c}$ and $\theta_{b}$ from the priors $\pi^{c}$ and $\pi^{b}$, respectively (lines $2-10$ of Algorithm 1). At iteration $r>1$, as done in standard SMC-ABC, we sample a particle $\theta_{c,r}^{(h)}=\theta_{c}$ for the continuous parameters from the weighted set $\{\Theta_{c,r-1},{w}_{r-1}\}$ (line 17 of Algorithm 1). It is then perturbed to $\theta_{c}^{*}$ with a Gaussian proposal kernel $\mathcal{K}^{c}_{r}$ (defined as in (8)) and proposed in [14] (line 18 of Algorithm 1). In particular, $\theta_{c}^{*}$ is a realization of $\mathcal{N}(\theta_{c},\hat{\Sigma}_{c,r})$, where $\hat{\Sigma}_{c,r}$ is twice the weighted covariance matrix obtained from the previous population $\{\Theta_{c,r-1},{w}_{r-1}\}$ of continuous parameters. Moreover, differently from the classical SMC-ABC, each entry $\theta_{b}^{u}$ of a particle $\theta_{b,r}^{(h)}=\theta_{b}$ for a binary parameter is drawn from a Bernoulli distribution with “success” probability given by the sample mean obtained from the respective particles of previous population $\Theta_{b,r-1}$ (line 19 of Algorithm 1). Such sampled value $\theta_{b}^{u}$, for $u=1,\ldots,b_{n}$, is then perturbed to $\theta_{b}^{*,u}$ with a Bernoulli-type kernel (line 20 of Algorithm 1). In particular, $\theta_{b}^{u}$ is either kept, i.e., $\theta_{b}^{*,{u}}=\theta_{b}^{u}$ (with probability $q_{\textrm{stay}}$) or perturbed to $\theta_{b}^{*,{u}}=1-\theta_{b}^{u}$ (with probability $1-q_{\textrm{stay}}$). An explicit expression of such kernel $\mathcal{K}_{r}^{b}$ (in the style of (8)) is given by
@@
11:Initialize the weights by setting each entry of $w_{1}=(w_{1}^{(1)},\ldots,w_{1}^{({M})})$ to $1/{M}$
@@ 3.3.1 Algorithm ingredients
$\begin{split}\textrm{Exp}(\Delta)&:=\begin{pmatrix}e^{-\Gamma\Delta}\left(\mathbb{I}_{3N}+\Gamma\Delta\right)&e^{-\Gamma\Delta}\Delta\\ -\Gamma^{2}e^{-\Gamma\Delta}\Delta&e^{-\Gamma\Delta}\left(\mathbb{I}_{3N}-\Gamma\Delta\right)\end{pmatrix}=:\begin{pmatrix}\vartheta(\Delta)&\kappa(\Delta)\\ \vartheta^{\prime}(\Delta)&\kappa^{\prime}(\Delta)\end{pmatrix},\\ \textrm{Cov}(\Delta)&:=\begin{pmatrix}\frac{1}{4}\Gamma^{-3}\Sigma^{2}\left(\mathbb{I}_{3N}+\kappa(\Delta)\vartheta^{\prime}(\Delta)-\vartheta^{2}(\Delta)\right)&\frac{1}{2}\Sigma^{2}\kappa^{2}(\Delta)\\ \frac{1}{2}\Sigma^{2}\kappa^{2}(\Delta)&\frac{1}{4}\Gamma^{-1}\Sigma^{2}\left(\mathbb{I}_{3N}+\kappa(\Delta)\vartheta^{\prime}(\Delta)-\kappa^{\prime 2}(\Delta)\right)\end{pmatrix}.\end{split}$ (10)
@@
Our proposed summary statistics of $y$ consist of functions that capture both the marginal behavior of each population $Y^{k}$, $k=1,\ldots,N$, and the interactions between populations $Y^{k}$ and $Y^{j}$, $j,k=1,\ldots,N$, $j\neq k$.
@@
[…]
@@
[…]
@@
$D(s(y),s(\tilde{y}_{\theta}))$ ::=:= $v_{1}\frac{1}{N}\sum\limits_{k=1}^{N}\textrm{IAE}(\hat{S}_{k},\tilde{S}_{k})+v_{2}\frac{1}{N}\sum\limits_{k=1}^{N}\textrm{IAE}(\hat{f}_{k},\tilde{f}_{k})$ (13)

### v4

@@ 1 Introduction
When embedded into a statistical inference procedure, this standard numerical scheme may therefore yield wrong estimation results, make the inference algorithm computationally infeasible or lead to ill-conditioned estimation methods [8, 13, 29, 31]. To obtain a reliable and efficient numerical simulation method, we further develop the structure-preserving splitting procedure proposed in [1] (see also [2, 40] for similar methods) to our $N$-population SDE model (second contribution). In contrast to the Euler-Maruyama scheme, which is based on truncating a stochastic Taylor series, the idea behind the splitting approach is to divide the unsolvable equation into solvable subequations, and to compose their solutions in a proper way [6, 25]. The constructed splitting scheme for the stochastic $N$-population neural mass model is based on a Hamiltonian type re-formulation of the SDE and successfully handles multiple interacting components, allowing to accurately simulate the complex dynamics of coupled neural populations.
@@
We denote by $0_{d}$ the $d$-dimensional zero vector, by $\mathbb{O}_{d}$ the $d\times d$-dimensional zero matrix, by $\mathbb{I}_{d}$ the $d\times d$-dimensional identity matrix, and by $\textrm{diag}[a_{1},\ldots,a_{d}]$ a $d\times d$-dimensional diagonal matrix with diagonal entries $a_{1},\ldots,a_{d}$. The transpose is denoted by and the Euclidean norm by $\left\lVert\cdot\right\rVert$. We sometimes omit the time index of a stochastic process and use, e.g., $(X(t))_{t\in[0,T]}$ and $X$ interchangeably.
@@ 2.1 One neural population
The meaning and typical values of the model parameters $A$, $B$, $a$, $b$, $C_{1}$, $C_{2}$, $C_{3}$, $C_{4}$, $\nu_{\max}$, $\gamma$ and $v_{0}$ are reported in Table 1 (see [1, 19] and the references therein). The average incoming membrane potential of the main cells, i.e., the process
$\vskip-5.69046pty(t):=x_{2}(t)-x_{3}(t),\vskip-5.69046pt$
@@
Table 1: Typical parameter values for the Jansen and Rit Neural Mass Model from the literature.
Parameter Meaning Typical value
AA Average excitatory synaptic gain $3.25$ mV
@@
$C_{1}$, $C_{2}$ Avg. no. of synaptic contacts in the excitatory feedback loop $C$, 0.8C0.8\ C
$C_{3}$, $C_{4}$ Avg. no. of synaptic contacts in the inhibitory feedback loop 0.25C0.25\ C, 0.25C0.25\ C
$\nu_{\max}$ Maximum firing rate (Maximum of the sigmoid function) 55 s-1
$v_{0}$ Value for which $50\%$ of the maximum firing rate is attained 66 mV
@@
An illustration of a simulated trace with parameters chosen to produce $\alpha$-waves (neural oscillations in the 8–12 Hz frequency band) is provided in Figure 1A. The model can also produce more complex behavior, such as brain signals occurring before and during epileptic seizures, by increasing the excitation-inhibition-ratio $A/B$ [44]. In Figure 1B, regular EEG activity is produced by using $\mu=90$, $\sigma=500$, $A=3.25$ and the typical values reported in Table 1. Increasing $A$ causes the model to generate sporadic and frequently occurring spikes ($A=3.5$ and $A=3.6$) and rhythmic discharge of spikes ($A=4.3$), see Figure 1C-1E.
Figure 1: Traces of one neural population with different types of activity. Simulated paths of the process $Y$ (4) of model (3). Panel A: Trace showing $\alpha$-rhythmic activity using $C=134.263$, $\mu=202.547$, $\sigma=1859.211$ (values taken from [8]), and $A=3.25$. Panels B-E: Traces describing activity occurring during epileptic seizures using $C=135$, $\mu=90$, $\sigma=500$ and different values of AA. Panel B: Regular EEG, $A=3.25$. Panel C: Sporadic spikes, $A=3.5$. Panel D: Frequently occurring spikes, $A=3.6$. Panel E: Rhythmic discharge of spikes, $A=4.3$. All remaining parameter values correspond to those listed in Table 1. The signals resemble experimental stereo EEG recordings (cf. Figure 3 in [44]).
@@ 2.2 Multiple coupled neural populations
$dX^{k}_{5}(t)=\Bigl{[}A_{k}a_{k}\Bigl{(}\mu_{k}+C_{2,k}\textrm{sig}\left(C_{1,k}X^{k}_{1}(t)\right)+\hskip-8.53581pt\sum\limits_{j=1,j\neq k}^{N}\rho_{jk}K_{jk}X_{1}^{j}(t)\Bigr{)}-2a_{k}X^{k}_{5}(t)-a_{k}^{2}X^{k}_{2}(t)\Bigr{]}dt+\sigma_{k}dW^{k}_{5}(t).$ (5)
@@
where $L>0$ is a coupling strength parameter and the parameter $0\ll c<1$ describes how fast the network coupling strength decreases with increasing distance between populations.
@@
The parameters of the stochastic multi-populatuion JR-NMM thus consist of continuous model parameters (cf. Table 1), two further continuous coupling strength parameters $L>0$ and $0\ll c<1$ (cf. (5) and (6)) and binary coupling direction (network) parameters $\rho_{jk}$ $1,\ldots,N, jkj\neq k$ (cf. (5)). Moreover, the observed component of the model is given by the $N$-dimensional process
@@
Similar to Section 2.1, we present a short simulation study where we generate paths of $Y$ (7), focusing on the newly introduced coupling term. According to [44], not only the excitation-inhibition-ratio $A/B$ is relevant for epileptic behavior, but also the coupling strengths and directions between neural groups play a crucial role. This is illustrated in Figure 2, where simulated activity of $N=4$ neural populations under different coupling regimes is shown. In the left panels, no coupling occurs, i.e., all $\rho_{jk}$-parameters are set to zero. In the middle and right panels, there is a unidirectional (cascade) coupling structure (illustrated in Figure 3a of Section 4), i.e, $\rho_{12}=\rho_{23}=\rho_{34}=1$, for different coupling strengths $L=K_{12}=K_{23}=K_{34}$. The activity of a passive site (no epileptic spikes occur without input from other populations) strongly depends on that of an active site (epileptic spikes occur without input from other populations). In Population $1$, we set $A_{1}=3.6$ to obtain spiking activity. In the remaining populations, the typical values of Table 1, $\mu=90$ and $\sigma=500$ are used. When there is no coupling, no activation of Populations 2–4 occurs (left panels). Introducing coupling and setting $L=K_{12}=K_{23}=K_{34}=300$ (central panels) leads to a dependence of Populations 2–4 on Population $1$. When the coupling is strong enough ($L=K_{12}=K_{23}=K_{34}=500$, right panels), rhythmic synchronization occurs. A similar behavior for two populations has also been observed for the original JR-NMM (cf. Figure 6 in [44]).
Figure 2: Cascade network of four populations with one active population. Simulated paths of $Y$ (7) for $N=4$ neural populations. The typical values of Table 1, $\mu=90$ and $\sigma=500$ are used, except for Population 11 where $A_{1}=3.6$ to activate it. Panel A: $\rho_{jk}=0$, $1,\ldots,4, jkj\neq k$. Panels B-C: $\rho_{12}=\rho_{23}=\rho_{34}=1$. The coupling strength parameters $K_{12}$, $K_{23}$ and $K_{34}$ equal $L=300$ and $L=500$ in Panel B and Panel C, respectively.
@@ 3.1 Standard ABC schemes
$\pi(\theta|y)\approx\pi_{\textrm{ABC}}^{\delta}(\theta|s(y))\propto\int\mathbbm{1}_{\{D(s(y),s(\tilde{y}_{\theta^{\prime}}))<\delta\}}\pi(\theta)\mathcal{L}(s(\tilde{y}_{\theta^{\prime}})|\theta)ds.$
@@
At iteration $r>1$, a particle $\theta$ is initially sampled from the set of kept candidates $\Theta_{r-1}$ of the previous iteration with the corresponding weights $w_{r-1}$ and then perturbed to a value $\theta^{*}\sim\mathcal{K}_{r}(\cdot|\theta)$, where $\mathcal{K}_{r}$ is a suitable perturbation kernel. For continuous parameters $\theta$, $\mathcal{K}_{r}$ is commonly assumed to be Gaussian (see, e.g., [12, 15, 35]), even though other possibilities have been proposed [28]. Here, we consider the multivariate Gaussian sampler as proposed in [15], that is a candidate $\theta^{*}$ is obtained from $\mathcal{N}(\theta,\hat{\Sigma}_{r})$, where $\hat{\Sigma}_{r}$ is twice the weighted empirical covariance matrix obtained from the previous population $\{\Theta_{r-1},w_{r-1}\}$ (cf. Appendix B.2 for an alternative Gaussian proposal sampler). Synthetic data $\tilde{y}_{\theta^{*}}$ are then simulated conditioned on the perturbed particle $\theta^{*}$, which is accepted if $d(s(y),s(\tilde{y}_{\theta^{*}}))<\delta_{r}$, with $\delta_{r}<\delta_{r-1}$. This is repeated until $M$ particles $\Theta_{r}=(\theta_{r}^{(1)},\ldots,\theta_{r}^{(M)})$ have been accepted. In particular, after the $h\text{-th}$ particle $\theta_{r}^{(h)}$, for $h\in\{1,\ldots,M\}$, has been accepted, the corresponding important weight $w_{r}^{(h)}$ is updated as
@@
This procedure is then repeated over several iterations (with decreasing threshold $\delta_{r}<\delta_{r-1}$ chosen as a certain percentile of the distances computed at the previous iteration r1r-1) until a suitable stopping criterion is reached (e.g., the acceptance rate of particles drops below a prefixed level or a certain number of model simulations, denoted by $n_{\textrm{sim}}$, has been reached). The final SMC-ABC posterior is obtained by sampling from the particles kept at the last iteration $r_{\textrm{last}}$ with probabilities given by the corresponding normalized weights. A detailed formulation of the standard SMC-ABC can be found, e.g., in Algorithm 2 of [28].
@@ 3.2 Adapted SMC-ABC algorithm for network inference: nSMC-ABC
At iteration $r=1$, we run the standard acceptance-rejection ABC, sampling parameter candidates $\theta_{c}$ and $\theta_{b}$ from the priors $\pi^{c}$ and $\pi^{b}$, respectively (lines 2102-10 of Algorithm 1). At iteration $r>1$, as done in standard SMC-ABC, we sample a particle $\theta_{c,r}^{(h)}=\theta_{c}$ for the continuous parameters from the weighted set $\{\Theta_{c,r-1},{w}_{r-1}\}$ (line 17 of Algorithm 1). It is then perturbed to $\theta_{c}^{*}$ with a Gaussian proposal kernel $\mathcal{K}^{c}_{r}$ (defined as in (8)) and proposed in [15] (line 18 of Algorithm 1). In particular, $\theta_{c}^{*}$ is a realization of $\mathcal{N}(\theta_{c},\hat{\Sigma}_{c,r})$, where $\hat{\Sigma}_{c,r}$ is twice the weighted covariance matrix obtained from the previous population $\{\Theta_{c,r-1},{w}_{r-1}\}$ of continuous parameters. Moreover, differently from the classical SMC-ABC, each entry $\theta_{b}^{u}$ of a particle $\theta_{b,r}^{(h)}=\theta_{b}$ for a binary parameter is drawn from a Bernoulli distribution with “success” probability given by the sample mean obtained from the respective particles of previous population $\Theta_{b,r-1}$ (line 19 of Algorithm 1). Such sampled value $\theta_{b}^{u}$, for $u=1,\ldots,b_{n}$, is then perturbed to $\theta_{b}^{*,u}$ with a Bernoulli-type kernel (line 20 of Algorithm 1). In particular, $\theta_{b}^{u}$ is either kept, i.e., $\theta_{b}^{*,{u}}=\theta_{b}^{u}$ (with probability $q_{\textrm{stay}}$) or perturbed to $\theta_{b}^{*,{u}}=1-\theta_{b}^{u}$ (with probability $1-q_{\textrm{stay}}$). An explicit expression of such kernel $\mathcal{K}_{r}^{b}$ (in the style of (8)) is given by
@@
11:Initialize the weights by setting each entry of $w_{1}=(w_{1}^{(1)},\ldots,w_{1}^{({M})})$ to 1/M1/{M}
@@ 3.3.1 Algorithm ingredients
[…]
@@
Our proposed summary statistics of $y$ consist of functions that capture both the marginal behavior of each population $Y^{k}$, $k=1,\ldots,N$, and the interactions between populations $Y^{k}$ and $Y^{j}$, $1,\ldots,N, jkj\neq k$.
@@
$S_{k}(\nu)=\mathcal{F}\{R_{k}\}(\nu)=\int\limits_{-\infty}^{\infty}R_{k}(\tau)e^{-i2\pi\nu\tau}\ d\tau,\quad k=1,\ldots,N,$
@@
The cross-spectral densities $S_{jk}(\nu)=\mathcal{F}\{R_{jk}\}(\nu)=\int\limits_{-\infty}^{\infty}R_{jk}(\tau)e^{-i2\pi\nu\tau}\ d\tau$, $1,\ldots,N, j<k,j<k$, could be other suitable summary statistics. However, these are symmetric (i.e., $S_{jk}(\nu)=S_{kj}(\nu)$), and thus cannot determine the direction of a connection. Indeed, our experiments show that adding those functions to the set of summaries does not yield an additional benefit.
@@
$D(s(y),s(\tilde{y}_{\theta}))$ $:=$ $v_{1}\frac{1}{N}\sum\limits_{k=1}^{N}\textrm{IAE}(\hat{S}_{k},\tilde{S}_{k})+v_{2}\frac{1}{N}\sum\limits_{k=1}^{N}\textrm{IAE}(\hat{f}_{k},\tilde{f}_{k})$ (13)